hector has 7/8 of a pound of leftover Halloween candy. He plans on giving it away to his friends. if he plans on giving each person 1/16 of a pound, how many people can he give to?
thanks.

Answer by addingup(3677)

1/16x = 7/8
x = 7/8*16/1 = 14
-------------------------------------
He gives out 1/16 to each. He has 7/8. How many 1/16 fit into 7/8? We first find a common denominator, this is easy since 8*2 = 16. Do the same with the top: 7*2 = 14. So now we have 14/16. Do you see your answer? You have 14 pieces to give away out of a total of 16, same answer we got solving with algebra.
Feel free to email if I can be of help.